# Dynamics and lifetime of geometric excitations in moiré systems

### v1

@@ I Introduction
In this paper, we explore the graviton-like excitations in the moiré systems such as twisted bilayer graphene (TBG), where FCI phases have been observed [22]. We extract the essential physics using a simplified ideal flat band (IFB) model, revealing that GMs are significantly weaker compared to the FQH case [58, 59, 60, 61, 62, 63]. Our numerical results indicate a rapid decay of the GM lifetime as the density of states (DOS) increases, suggesting that the observed behavior may be strongly influenced by finite-size effects. In particular, with Coulomb interactions, a peak in the spectral function emerges in small systems just above the continuum boundary. However, in the thermodynamic limit where the DOS diverges in the continuum, this resonance peak may disappear. Additionally, we show analytically that the suppression of GM lifetimes is due to the strong scattering between different angular momentum sectors, despite the surprising robustness of guiding center rotational invariance of the ground states. Both our analytic and numerical results should be applicable for generic Chern bands, and we suggest a potential approach to stabilize GMs in experimental settings in such systems.
@@ II Moiré ideal flat bands and geometric excitations
$\mathcal{F}(\boldsymbol{k}_{1},\boldsymbol{k}_{2}):$ $=\braket{\boldsymbol{k}_{1}}{e^{i(\boldsymbol{k}_{1}-\boldsymbol{k}_{2})\cdot\boldsymbol{r}}}{\boldsymbol{k}_{2}}$
$=\mathcal{N}_{\boldsymbol{k}_{1}}\mathcal{N}_{\boldsymbol{k}_{2}}\sum_{\boldsymbol{b}}w_{\boldsymbol{b}}f^{\boldsymbol{k}_{1},\boldsymbol{k}_{2}}_{-\boldsymbol{b}},$ (5)
@@
$f^{\boldsymbol{k}_{1},\boldsymbol{k}_{2}}_{-\boldsymbol{b}}=\braket{\boldsymbol{k}_{1}}{e^{i(\boldsymbol{k}_{1}-\boldsymbol{k}_{2}+\boldsymbol{b})\cdot\boldsymbol{r}}}{\boldsymbol{k}_{2}}_{\text{LLL}}.$
@@
$H_{\text{int}}$ $=\sum_{\boldsymbol{k}_{1},\boldsymbol{k}_{2}\in\text{ BZ}}\sum_{\boldsymbol{q}}V(\boldsymbol{q})\mathcal{F}(\boldsymbol{k}_{1},\boldsymbol{k}_{1}+\boldsymbol{q})\mathcal{F}(\boldsymbol{k}_{2},\boldsymbol{k}_{2}-\boldsymbol{q})$
@@
$\hskip 28.45274ptw_{\boldsymbol{b}_{1}}w_{\boldsymbol{b}_{2}}f^{\boldsymbol{k}_{1},\boldsymbol{k}_{1}+\boldsymbol{q}}_{-\boldsymbol{b}_{1}}f^{\boldsymbol{k}_{2},\boldsymbol{k}_{2}-\boldsymbol{q}}_{-\boldsymbol{b}_{2}}c_{\boldsymbol{k}_{1}}^{\dagger}c_{\boldsymbol{k}_{2}}^{\dagger}c_{\boldsymbol{k}_{2}-\boldsymbol{q}}c_{\boldsymbol{k}_{1}+\boldsymbol{q}}.$ (6)
@@
because for translationally invariant ground states, the leading contribution of $\delta\hat{\bar{\rho}}_{\boldsymbol{q}\rightarrow 0}$ acts as the generator of area-preserving deformations [35, 36, 37, 15, 74]. Here $S_{\boldsymbol{q}}$ is the projected static structure factor and $\delta\hat{\bar{\rho}}_{\boldsymbol{q}}=\hat{\bar{\rho}}_{\boldsymbol{q}}-\left\langle\psi_{0}\right|\hat{\bar{\rho}}_{\boldsymbol{q}}\left|\psi_{0}\right\rangle$ is the regularized density operator, both projected to a single Chern band. In particular, for LLs $\hat{\bar{\rho}}_{\boldsymbol{q}}$ represents the guiding center density operator. On compact manifolds such as spheres or tori, however, the momentum $\boldsymbol{q}$ becomes a discrete quantum number, and accessing the long-wavelength limit requires taking the thermodynamic limit.
@@
where $\left(q_{x}\pm iq_{y}\right)^{2}$ represents a chiral d-wave symmetry (the magnetic length $\ell_{B}\equiv 1$) and the LLL form factor $e^{-|\boldsymbol{q}|^{2}/2}$ has been absorbed into $V({\boldsymbol{q}})$. The chiral graviton operator $\hat{O}_{\pm}$ and the SMA operator $\delta\hat{\bar{\rho}}_{\boldsymbol{q}\rightarrow 0}$ in gapped FQH states are related to each other through the LLL Ward identities  [^1] The detailed discussion was given in the Supplement Material of Ref [44]. These two approaches differ in their universality and sensitivity to interaction details: The SMA operator is universal, as it captures neutral excitations arising from density modulations of the ground state. In contrast, chiral graviton operators are influenced by specific interaction details and characterize global metric deformations. Both the SMA operator and the chiral graviton operators are experimentally accessible through Raman scattering [30, 80, 39, 45, 78, 11], their measured spectral functions are related to each other by sum rules [39, 42, 45]. Chiral graviton operators can also be directly realized by tilting the magnetic field [81, 82, 76, 38] or applying acoustic waves [75]. Both operators can be extended to IFBs by appropriately adjusting the form factors to account for the discrete translational symmetry of the Hamiltonian and the nonuniform quantum geometry.
@@ III Graviton lifetime in moiré IFBs
Figure 1: Spectral functions of the chiral graviton operator at the filling $\nu=1/3$ in moiré IFB and the LLL. All the spectral peaks shown in this paper have been normalized. The system sizes are distinguished by different colors. Arrows denote the boundary positions of the excitation continuum. The energy scale is determined by the Coulomb energy $E_{C}=e^{2}/(4\pi\epsilon l)$, where the characteristic length scale $l=\ell_{B}$ for LLL and $l=\sqrt{\sqrt{3}/4\pi}a_{M}$ for IFB. Here $a_{M}$ is the moiré lattice constant. (a, b) In the IFB of cTBG, a sharp peak persists under $V_{C}$ even as the system size increases. However, with the $V_{1}$ pseudopotential, the energy doubles compared to the LLL and dives deeper into the continuum, while the peaks become significantly suppressed as the system size slightly grows. (c, d) In the LLL, clear and sharp peaks are observed across all system sizes for both the Coulomb interaction $V_{C}$ (dashed stems) and the $V_{1}$ pseudopotential (solid stems). The model Hamiltonian exhibits a particularly pronounced peak, consistent with previous studies on the LLL in torus geometry. These indicate a much shorter GM lifetime in moiré IFBs compared to LLs.
@@ V The moiré system as perturbed LLs
where $\boldsymbol{r}=\boldsymbol{r}_{1}-\boldsymbol{r}_{2}$ is the relative coordinate and $\boldsymbol{R}=\frac{\boldsymbol{r}_{1}+\boldsymbol{r}_{2}}{2}$ is the centre-of-mass coordinate, and the summation is over the smallest reciprocal lattice vectors (sRLV). One can prove analytically that, with the shortest range interaction for fermions, $V(\boldsymbol{r})=\nabla^{\prime\prime}\delta\left(\boldsymbol{r}\right)$, the isotropic Laughlin state at $\nu=1/3$ is the exact zero energy state for the effective LLL model  [^3] Detailed derivations are given in the supplementary material.. This is despite the explicit breaking of the continuous rotational symmetry of the Hamiltonian, which is evident in the anisotropic dependence of $\delta V(\boldsymbol{r}_{1},\boldsymbol{r}_{2})$ on the center-of-mass coordinate $\boldsymbol{R}$. We therefore see that the excited states will generally be affected by scattering that changes the total angular momentum of a pair of electrons, while the ground state remains isotropic.
@@
where in $\tilde{V}_{\boldsymbol{q},\boldsymbol{s},\boldsymbol{t}}=V(\boldsymbol{q})\left(1+\varepsilon_{\boldsymbol{q},\boldsymbol{s},\boldsymbol{t}}\right)$, $\boldsymbol{q}$ is the momentum transfer (Here we take $w_{0}=1$ without loss of generality). The broken continuous translational symmetry from the lattice system is encoded in the perturbation:
@@
Interestingly, even beyond the linearized special case in Eq.(12), the ground state wavefunctions are invariant as exact zero energy Laughlin states with arbitrary choices of $w_{\boldsymbol{b}_{i}}$ for $V(\boldsymbol{q})=V_{1}$, with the exact three-fold degeneracy at $\nu=1/3$. This implies that the universal properties of the GM still hold, including the chirality [77, 78, 43]. This observation has been confirmed by extensive numerical computations. We conjecture the universal behavior can be analytically proven by expanding Eq. 15 in the orthonormal basis of the generalized pseudopotentials (PPs). Crucially, the perturbation for nonzero $w_{\boldsymbol{b}_{i}}$ is holomorphic in $q$, apart from in $V_{1}(\boldsymbol{q})$, reflecting the chiral nature of the IFB. These holomorphic terms do not compromise the Hermiticity of the Hamiltonian in the torus geometry  [^4] As long as the inversion symmetry in $\boldsymbol{b}_{i}$ vectors is preserved, the Hamiltonian will remain Hermitian. Refer to the supplementary material for more details.. Because of this holomorphicity, the expansion in terms of generalized pseudopotentials only involves the $V^{+}_{1,m}$ terms  [^5] Detailed derivations are given in the supplementary material.:
@@
Figure 4: (a) Spectral function of $\hat{H}_{s}$ with different perturbation strengths. A Gaussian fit is applied to track the evolution of the GM peak, which diminishes rapidly with increasing perturbation strength. The $w_{1}$ in the legend is relative to the cTBG value. The spectral peaks decrease monotonically with larger $w_{1}$. (b) Spectral functions with modified Coulomb interaction $V_{\text{ZDS}}$ in tMoTe2. The system size is with $24$ orbitals at the hole filling $\nu=-2/3$. Incorporating the finite-thickness effect described by $V_{\text{ZDS}}$ shifts the GM peak further away from the continuum. This separation isolates the GM from other excitations and enhances its lifetime, as shown by the higher peaks. (c) Spectrum of $V_{\text{ZDS}}$ at $\lambda=0.2$. The red dot shows the position of the GM peak presented in b, gapped from the continuum. Here $(K_{1},K_{2})$ gives the position in the first Brillouin zone with $N_{1}=4$ as the number of sites along one of the primitive lattice vectors.
@@
The emergent rotational symmetry should also be robust in realistic systems or more general models (e.g., beyond the model interaction or the chiral limit), as long as the topological phase itself is robust with a large incompressible gap compared to the perturbations to the chiral interaction. This is because the perturbations to the chiral interaction are suppressed by the incompressibility gap, rendering the ground state (and thus the GM) only slightly anisotropic. In contrast, $\varepsilon_{\boldsymbol{q},\boldsymbol{s},\boldsymbol{t}}$ significantly perturbs the continuous gapped excitations of $\hat{H}_{LLL}$, scattering the excitations between different angular momentum sectors. In particular, the spin-2 modes present in the LLL spectrum, responsible for the narrow spectral peak, will scatter with other higher spin modes by $\varepsilon_{\boldsymbol{q},\boldsymbol{s},\boldsymbol{t}}$. This explanation is numerically confirmed by Fig. 4a, where the GM peak of $\hat{H}_{s}$ rapidly decreases with increasing $w_{1}$.
@@
It is useful to recall that microscopic Hamiltonians of the moiré systems (or lattice systems in general) only have discrete rotational symmetry (e.g. $C_{3}$ symmetry in TBG), one thus naively would expect the quantum fluid to have only discrete rotational symmetry both below or above the incompressibility gap. Geometric excitations from the ground state thus seem complicated, and not much can be said about their dynamics other than from numerical computations with finite system sizes. The results from numerical computations are challenging to interpret due to the small DOS, as discussed in the previous section. Generically one would argue that the GMs (which may not even be well defined due to discrete symmetry) are naturally weak and so the spectral peaks found in numerics are surprising. The discovery that the ground state (or any states below the gap) and thus the geometric excitation have emergent continuous rotational symmetry, while gapped excitations do not, leads to two main messages: firstly it gives a firm understanding on why GMs in the continuum will have vanishing lifetime despite the numerical results from small systems; secondly the GMs are still well defined and they can in principle be measured when properly tuned, as we will discuss in detail in the next section.

### v2

@@ I Introduction
Figure 1: Schematic illustrating the shorter lifetime of graviton modes (GMs) in moiré Chern bands compared to Landau levels (LLs). In both systems, the GMs $\psi_{G}$ are excited by incoming polarized beams (red) from the same topological ground state $\psi_{0}$. These modes possess well-defined “spins” due to an emergent guiding-center rotational symmetry. In LLs, all gapped excitations $\psi_{i}$ maintain continuous rotational symmetry, ensuring that spin-2 graviton modes can only scatter into sectors with the same spin, depicted by the quadrupole structure in the figure. In contrast, moiré Chern bands lack continuous rotational symmetry due to the underlying lattice structure. Changing from the LLs to moiré Chern bands can be considered as adding perturbations that have little effect on the ground state and GM, but completely reorganize the excited states from $\{\psi_{i}\}$ to $\{\tilde{\psi}_{i}\}$. Consequently, the highly dense excitations $\tilde{\psi}_{i}$ do not possess well-defined spins, leading to graviton modes scattering into multiple channels. As a result, in moiré systems, there are no resonance peaks from the outgoing polarized beam (blue), implying a significantly short GM lifetime, whereas in LLs, a sharp spectral peak can be observed when scanning different frequencies.
@@
In this paper, we explore the graviton-like excitations in the moiré systems such as twisted bilayer graphene (TBG), where FCI phases have been observed [22]. We extract the essential physics using a simplified ideal flat band (IFB) model, revealing that GMs are significantly weaker compared to the FQH case [58, 59, 60, 61, 62, 63]. Our numerical results indicate a rapid decay of the GM lifetime as the density of states (DOS) increases, suggesting that the observed behavior may be strongly influenced by finite-size effects. In particular, with Coulomb interactions, a peak in the spectral function emerges in small systems just above the continuum boundary. However, in the thermodynamic limit where the DOS diverges in the continuum, this resonance peak may disappear. Additionally, we show analytically that the suppression of GM lifetimes is due to the strong scattering between different angular momentum sectors, despite the surprising robustness of guiding center rotational invariance of the ground states, as illustrated in Fig. 1. Both our analytic and numerical results should be applicable for generic Chern bands, and we suggest a potential approach to stabilize GMs in experimental settings in such systems.
@@ II Moiré ideal flat bands and geometric excitations
$\mathcal{F}(\boldsymbol{k}_{1},\boldsymbol{k}_{2}):$ $=\braket{\boldsymbol{k}_{1}}{e^{i(\boldsymbol{k}_{1}-\boldsymbol{k}_{2})\cdot\boldsymbol{r}}}{\boldsymbol{k}_{2}}$ (5)
$=\mathcal{N}_{\boldsymbol{k}_{1}}\mathcal{N}_{\boldsymbol{k}_{2}}\sum_{\boldsymbol{b}}w_{\boldsymbol{b}}f^{\boldsymbol{k}_{1},\boldsymbol{k}_{2}}_{-\boldsymbol{b}},$
@@
$f^{\boldsymbol{k}_{1},\boldsymbol{k}_{2}}_{-\boldsymbol{b}}=\braket{\boldsymbol{k}_{1}}{e^{i(\boldsymbol{k}_{1}-\boldsymbol{k}_{2}+\boldsymbol{b})\cdot\boldsymbol{r}}}{\boldsymbol{k}_{2}}_{\text{LLL}}.$ (6)
@@
$H_{\text{int}}$ $=\sum_{\boldsymbol{k}_{1},\boldsymbol{k}_{2}\in\text{ BZ}}\sum_{\boldsymbol{q}}V(\boldsymbol{q})\mathcal{F}(\boldsymbol{k}_{1},\boldsymbol{k}_{1}+\boldsymbol{q})\mathcal{F}(\boldsymbol{k}_{2},\boldsymbol{k}_{2}-\boldsymbol{q})$ (7)
@@
$\hskip 28.45274ptw_{\boldsymbol{b}_{1}}w_{\boldsymbol{b}_{2}}f^{\boldsymbol{k}_{1},\boldsymbol{k}_{1}+\boldsymbol{q}}_{-\boldsymbol{b}_{1}}f^{\boldsymbol{k}_{2},\boldsymbol{k}_{2}-\boldsymbol{q}}_{-\boldsymbol{b}_{2}}c_{\boldsymbol{k}_{1}}^{\dagger}c_{\boldsymbol{k}_{2}}^{\dagger}c_{\boldsymbol{k}_{2}-\boldsymbol{q}}c_{\boldsymbol{k}_{1}+\boldsymbol{q}}.$
@@
because for translationally invariant ground states, the leading contribution of $\delta\hat{\bar{\rho}}_{\boldsymbol{q}\rightarrow 0}$ acts as the generator of area-preserving deformations [35, 36, 37, 15, 74, 75, 76]. Here $S_{\boldsymbol{q}}$ is the projected static structure factor and $\delta\hat{\bar{\rho}}_{\boldsymbol{q}}=\hat{\bar{\rho}}_{\boldsymbol{q}}-\left\langle\psi_{0}\right|\hat{\bar{\rho}}_{\boldsymbol{q}}\left|\psi_{0}\right\rangle$ is the regularized density operator, both projected to a single Chern band. In particular, for LLs $\hat{\bar{\rho}}_{\boldsymbol{q}}$ represents the guiding center density operator. On compact manifolds such as spheres or tori, however, the momentum $\boldsymbol{q}$ becomes a discrete quantum number, and accessing the long-wavelength limit requires taking the thermodynamic limit.
@@
where $\left(q_{x}\pm iq_{y}\right)^{2}$ represents a chiral d-wave symmetry (the magnetic length $\ell_{B}\equiv 1$) and the LLL form factor $e^{-|\boldsymbol{q}|^{2}/2}$ has been absorbed into $V({\boldsymbol{q}})$. The chiral graviton operator $\hat{O}_{\pm}$ and the SMA operator $\delta\hat{\bar{\rho}}_{\boldsymbol{q}\rightarrow 0}$ in gapped FQH states are related to each other through the LLL Ward identities  [^1] The detailed discussion was given in the Supplement Material of Ref [44]. These two approaches differ in their universality and sensitivity to interaction details: The SMA operator is universal, as it captures neutral excitations arising from density modulations of the ground state. In contrast, chiral graviton operators are influenced by specific interaction details and characterize global metric deformations. Both the SMA operator and the chiral graviton operators are experimentally accessible through Raman scattering [30, 82, 39, 45, 80, 11], and their measured spectral functions are related to each other by sum rules [39, 42, 45]. Chiral graviton operators can also be directly realized by tilting the magnetic field [83, 84, 78, 38] or applying acoustic waves [77]. Both operators can be extended to IFBs by appropriately adjusting the form factors to account for the discrete translational symmetry of the Hamiltonian and the nonuniform quantum geometry.
@@ III Graviton lifetime in moiré IFBs
Figure 2: Spectral functions of the chiral graviton operator at the filling $\nu=1/3$ in moiré IFB and the LLL. All the spectral peaks shown in this paper have been normalized. The system sizes are distinguished by different colors. Arrows denote the boundary positions of the excitation continuum. The energy scale is determined by the Coulomb energy $E_{C}=e^{2}/(4\pi\epsilon l)$, where the characteristic length scale $l=\ell_{B}$ for LLL and $l=\sqrt{\sqrt{3}/4\pi}a_{M}$ for IFB. Here $a_{M}$ is the moiré lattice constant. (a, b) In the IFB of cTBG, a sharp peak persists under $V_{C}$ even as the system size increases. However, with the $V_{1}$ pseudopotential, the energy doubles compared to the LLL and dives deeper into the continuum, while the peaks become significantly suppressed as the system size slightly grows. (c, d) In the LLL, clear and sharp peaks are observed across all system sizes for both the Coulomb interaction $V_{C}$ (dashed stems) and the $V_{1}$ pseudopotential (solid stems). The model Hamiltonian exhibits a particularly pronounced peak, consistent with previous studies on the LLL in torus geometry. These indicate a much shorter GM lifetime in moiré IFBs compared to LLs. A Gaussian fit is applied to highlight the GM peak in the plots visually.
@@ V The moiré system as perturbed LLs
where $\boldsymbol{r}=\boldsymbol{r}_{1}-\boldsymbol{r}_{2}$ is the relative coordinate and $\boldsymbol{R}=\frac{\boldsymbol{r}_{1}+\boldsymbol{r}_{2}}{2}$ is the centre-of-mass coordinate, and the summation is over the smallest reciprocal lattice vectors (sRLV). One can prove analytically that, with the shortest range interaction for fermions, $V(\boldsymbol{r})=\nabla^{\prime\prime}\delta\left(\boldsymbol{r}\right)$, the isotropic Laughlin state at $\nu=1/3$ is the exact zero energy state for the effective LLL model. This is despite the explicit breaking of the continuous rotational symmetry of the Hamiltonian, which is evident in the anisotropic dependence of $\delta V(\boldsymbol{r}_{1},\boldsymbol{r}_{2})$ on the center-of-mass coordinate $\boldsymbol{R}$. We therefore see that the excited states will generally be affected by scattering that changes the total angular momentum of a pair of electrons, while the ground state remains isotropic.
@@
where in $\tilde{V}_{\boldsymbol{q},\boldsymbol{s},\boldsymbol{t}}=V(\boldsymbol{q})\left(1+\varepsilon_{\boldsymbol{q},\boldsymbol{s},\boldsymbol{t}}\right)$, $\boldsymbol{q}$ is the momentum transfer (Here we take $w_{0}=1$ without loss of generality). By replacing the IFB creation and annihilation operators with LLL operators, we can map the IFB system to an LLL problem with the same representation for the Hamiltonian in momentum basis. The broken continuous translational symmetry from the original lattice system is encoded in the perturbation:
@@
Interestingly, even beyond the linearized special case in Eq.(13), the ground state wavefunctions are invariant as exact zero energy Laughlin states with arbitrary choices of $w_{\boldsymbol{b}_{i}}$ for $V(\boldsymbol{q})=V_{1}$, with the exact three-fold degeneracy at $\nu=1/3$. This implies that the universal properties of the GM still hold, including the chirality [79, 80, 43], which has been confirmed by extensive numerical computations. The universal behavior can be analytically proven by expanding Eq. 16 in the orthonormal basis of the generalized pseudopotentials (PPs). Crucially, the perturbation for nonzero $w_{\boldsymbol{b}_{i}}$ is holomorphic in $q$, apart from in $V_{1}(\boldsymbol{q})$, reflecting the chiral nature of the IFB. These holomorphic terms do not compromise the Hermiticity of the Hamiltonian in the torus geometry  [^3] As long as the inversion symmetry in $\boldsymbol{b}_{i}$ vectors is preserved, the Hamiltonian will remain Hermitian. Refer to the supplementary material for more details.. Because of this holomorphicity, the expansion in terms of generalized pseudopotentials only involves the $V^{+}_{1,m}$ terms  [^4] Detailed derivations are given in the supplementary material.:
@@
Figure 5: Spectral function of $\hat{H}_{s}$ with different perturbation strengths. The GM peak diminishes rapidly with increasing perturbation strength. The $w_{1}$ in the legend is relative to the cTBG value. The spectral peaks decrease monotonically with larger $w_{1}$.
@@
The emergent rotational symmetry should also be robust in realistic systems or more general models (e.g., beyond the model interaction or the chiral limit), as long as the topological phase itself is robust with a large incompressible gap compared to the perturbations to the chiral interaction. This is because the perturbations to the chiral interaction are suppressed by the incompressibility gap, rendering the ground state (and thus the GM) only slightly anisotropic. In contrast, $\varepsilon_{\boldsymbol{q},\boldsymbol{s},\boldsymbol{t}}$ significantly perturbs the continuous gapped excitations of $\hat{H}_{LLL}$, scattering the excitations between different angular momentum sectors. In particular, the spin-2 modes present in the LLL spectrum, responsible for the narrow spectral peak, will scatter with other higher spin modes by $\varepsilon_{\boldsymbol{q},\boldsymbol{s},\boldsymbol{t}}$. This explanation is numerically confirmed by Fig. 5, where the GM peak of $\hat{H}_{s}$ rapidly decreases with increasing $w_{1}$.
@@
It is useful to recall that microscopic Hamiltonians of the moiré systems (or lattice systems in general) only have discrete rotational symmetry (e.g., $C_{3}$ symmetry in TBG), one thus naively would expect the quantum fluid to have only discrete rotational symmetry both below or above the incompressibility gap. Geometric excitations from the ground state thus seem complicated, and not much can be said about their dynamics other than from numerical computations with finite system sizes. The results from numerical computations are challenging to interpret due to the small DOS, as discussed in the previous section. Generically one would argue that the GMs (which may not even be well defined due to discrete symmetry) are naturally weak and so the spectral peaks found in numerics are surprising. The discovery that the ground state (or any states below the gap) and thus the geometric excitation have emergent continuous rotational symmetry, while gapped excitations do not, leads to two main messages: firstly it gives a firm understanding on why GMs in the continuum will have vanishing lifetime despite the numerical results from small systems; secondly the GMs are still well defined and they can in principle be measured when properly tuned, as we will discuss in detail in the next section.